### v2

# Learning Non-Ideal Single Vortex Flows Using the Differentiable Vortex Particle Method

Abstract
This study extends the differentiable vortex particle method (DVPM) beyond idealized flow scenarios to encompass more realistic, non-ideal conditions, including viscous flow and flow subjected to non-conservative body forces. We establish the Lamb-Oseen vortex as a benchmark case, representing a fundamental viscous single vortex flow in fluid mechanics. This selection offers significant analytical advantages, as the Lamb-Oseen vortex possesses an exact analytical solution derived from the Navier-Stokes (NS) equations, thereby providing definitive ground truth data for training and validation purposes. Through rigorous evaluation across a spectrum of Reynolds numbers, we demonstrate that DVPM achieves superior accuracy in modeling the Lamb-Oseen vortex compared to conventional convolutional neural networks (CNNs) and physics-informed neural networks (PINNs). Our results substantiate DVPM’s robust capabilities in modeling non-ideal single vortex flows, establishing its distinct advantages over contemporary deep learning methodologies in fluid dynamics applications. The dataset and source code are publicly available on GitHub at the following link: https://github.com/jh36714753/Learning_Non-Ideal_Single_Vortex_Flows.
1 Introduction
Vortex flows represent a fundamental class of fluid phenomena prevalent throughout nature. They manifest in diverse forms, such as turbulent structures, swirling water patterns, turbomachinery flows, and atmospheric systems, including hurricanes. Despite their ubiquity and significant importance, the complete understanding of vortex flow dynamics remains an open challenge in contemporary scientific research.
The concept of hidden fluid mechanics was introduced by Raissi et al. [8] in 2020, referring to fluid mechanical phenomena not explicitly represented by governing equations but instead embedded within flow visualizations. This paradigm aims to decode and learn fluid dynamics directly from visualization data. Currently, there are three principal machine learning approaches to hidden fluid mechanics: the PINNs-based method [8], the CNN-based method [15], and the DVPM-based approach [2]. In their pioneering PINNs-based method, Raissi et al. [8] enhanced the conventional PINNs framework by incorporating an additional scalar variable representing marked material concentration to visualize flow patterns. This methodology demonstrated remarkable success in addressing various physical and biomedical applications. Subsequently, Zhang et al. [15] developed the CNN-based approach, which integrates physical constraints into the loss function of the CNN framework. Their method effectively utilized particle image velocimetry (PIV) data for flow prediction. More recently, Deng et al. [2] proposed the DVPM-based method, introducing a deep learning-based vortex particle method specifically optimized for vortex flows. This specialized approach has exhibited superior performance to PINNs and CNN methods in analyzing vortical flow structures.
The DVPM, introduced by Deng et al. [2], which integrates deep learning techniques with traditional fluid dynamics approaches. This innovative method harnesses the differentiable properties of neural networks to enhance vortex flow prediction and analysis. However, the method’s initial implementation exhibits several limitations that constrain its practical applicability. Specifically, it is restricted to highly simplified vortex flow scenarios, including two-dimensional flows, inviscid conditions, and cases without non-conservative body forces. These constraints significantly limit its extension to more complex, real-world vortex flows. Additionally, a critical consideration arises regarding the method’s training dataset, which partially relies on traditional vortex particle method simulations. From the perspective of fluid mechanics researchers, these simulations may not provide sufficient accuracy to serve as reliable ground truth data.
This study extends the differentiable vortex particle method to encompass more general, non-ideal conditions, including viscous flow and flow influenced by non-conservative body forces. As a benchmark case, we introduce the viscous single vortex flow, specifically the Lamb-Oseen vortex. The Lamb-Oseen vortex, with its analytical solution derived directly from the NS equations, provides ground truth data for model training and validation. Through rigorous evaluation across various Reynolds numbers, we demonstrate that the differentiable vortex particle method achieves superior accuracy compared to the CNN-based and PINN-based approaches. Additionally, we validate our model’s capability to accurately predict single vortex flow dynamics under the combined influence of viscosity and non-conservative body forces. These results substantiate the robust performance of the differentiable vortex particle method in vortex flow modeling and establish its significant advantages over contemporary deep learning techniques.
2 Related work
The vortex method is a prominent computational technique that has garnered significant attention across multiple domains. Its applications extend beyond computational fluid dynamics (CFD) [4, 5, 6, 13] to computer graphics (CG) [9, 7, 12, 1, 3, 14, 10, 16, 11]. This methodology demonstrates the inherent advantages of meshless approaches, employs a Lagrangian viewpoint, and exhibits superior computational efficiency.
3 Physical model
For incompressible fluid flows, the vorticity transport equation, derived rigorously from the NS equations, establishes a comprehensive mathematical framework for characterizing the spatiotemporal evolution of vorticity in swirling flow fields:
$\frac{D\boldsymbol{\omega}}{Dt}=\frac{\partial\boldsymbol{\omega}}{\partial t}+\boldsymbol{u}\cdot\nabla\boldsymbol{\omega}=\boldsymbol{\omega}\cdot\nabla\boldsymbol{u}+\nu\nabla^{2}\boldsymbol{\omega}+\nabla\times\boldsymbol{b},$ (1)
where $\boldsymbol{\omega}$ represents the vorticity vector, $t$ signifies time, $\boldsymbol{u}$ denotes the velocity field, $\boldsymbol{b}$ corresponds to the non-conservative body force, and $\nu$ refers to the kinematic viscosity. For an inviscid, two-dimensional flow in the absence of non-conservative body forces, the equation reduces to $D\boldsymbol{\omega}/Dt=\mathbf{0}$, which explicitly demonstrates the Lagrangian conservation of vorticity (i.e., the vorticity of a fluid particle remains constant as it is advected along its trajectory). However, practical flows rarely satisfy these idealized conditions, leading to more complex flow phenomena characterized by intricate vortical dynamics.
Through the application of the Biot-Savart Law, we derive the vortex-induced velocity formulation:
[…]
where $K\left(\boldsymbol{x}-\boldsymbol{x}^{\prime}\right)$ denotes the kernel function, which is dependent on the displacement vector between the vorticity location and the point of interest in the flow field. In essence, Equation 2 indicates that the velocity at any point within the flow field can be determined by integrating the contributions of vorticity distributed throughout the domain, mediated by the kernel function.
Figure 1 presents the physical model employed in our computational framework. Within a two-dimensional flow field comprising $n$ discrete vortices, each vortex is characterized by three fundamental parameters: position $\boldsymbol{x}_{i}\prime$, vorticity $\omega_{i}$, and radius $\delta_{i}$ (assuming circular vortex geometry) for the i-th vortex. These vortices utilize a common kernel function $K\left(\frac{\boldsymbol{x}-\boldsymbol{x}^{\prime}}{\delta}\right)$. Subsequently, the velocity at any arbitrary position $\boldsymbol{x}$ within the flow field can be determined through:
$\boldsymbol{u}(\boldsymbol{x})=\sum_{i=1}^{n}K\left(\frac{\left|\boldsymbol{x}-\boldsymbol{x}_{i}^{\prime}\right|}{\delta_{i}}\right)\omega\left(\boldsymbol{x}_{i}^{\prime}\right)$ (3)
By Equation 3, estimating velocity throughout the flow fields necessitates knowledge of three key attributes of all vortices, along with their kernel function. To address this computational challenge, we implement a framework comprising four neural networks, each dedicated to predicting one of these essential physical quantities. In the two-dimensional flows, the temporal evolution of positional coordinates for n vortices can be expressed as $\left[\left(\boldsymbol{x}_{1}\right)_{t},\ldots,\left(\boldsymbol{x}_{n}\right)_{t}\right]$, while their vorticity distributions are represented by $\left[\left(\omega_{1}\right)_{t},\ldots,\left(\omega_{n}\right)_{t}\right]$, and their radial parameters are denoted by $\left[\left(\delta_{1}\right)_{t},\ldots,\left(\delta_{n}\right)_{t}\right]$. Consequently, we employ four distinct neural networks to characterize these fundamental physical quantities:
$\left\{\begin{array}[]{l}N_{\boldsymbol{x}}(t)=\left[\left(\boldsymbol{x}_{1}\right)_{t},\cdots,\left(\boldsymbol{x}_{n}\right)_{t}\right]\\ N_{\omega}(t)=\left[\left(\omega_{1}\right)_{t},\cdots,\left(\omega_{n}\right)_{t}\right]\\ N_{\delta^{2}}(t)=\left[\left(\delta_{1}^{2}\right)_{t},\cdots,\left(\delta_{n}^{2}\right)_{t}\right]\\ N_{K}\left(\frac{\left|\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\prime}\right|}{\delta_{i}}\right)=K\left(\frac{\left|\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\prime}\right|}{\delta_{i}}\right)\end{array}\right.$ (4)
Figure 1: The physical model in our framework. $K\left(\left|\boldsymbol{\boldsymbol{x}}-\boldsymbol{x}\prime\right|/\delta\right)$ represents the kernel function, whose value is determined by multiple parameters: the dependent variable, the Euclidean distance between the vortex core and the point of interest in the flow field $\left|\boldsymbol{x}-\boldsymbol{x}_{p}\right|$, and the vortex radius $\delta$. Each vortex is characterized by a distinct vorticity $\omega$, which reduces to a scalar quantity in two-dimensional flow configurations.
4 Method
Figure 2 presents a comprehensive overview of our framework. By analyzing historical RGB images, we can extract and track the temporal evolution of vortex attributes (including position, vorticity, and radius) in the Lagrangian frame of reference. We compute the Eulerian velocity field distribution by implementing the Biot-Savart law with an optimized kernel function. Subsequently, we obtain the RGB values for the subsequent time step by advecting the current image data through the calculated velocity field. For future state prediction, while the underlying physical model governs the vortex dynamics, the remaining computational pipeline remains consistent with the past analysis procedure.
Figure 2: An overview of our framework.
We formalize the vortex representation as $V_{t}=\left[\left(\boldsymbol{x}_{1},\omega_{1},\delta_{1}\right),\ldots,\left(\boldsymbol{x}_{n},\omega_{n},\delta_{n}\right)\right]_{t}$. In this framework, $V_{t}=T(t)$, where $T$ denotes the temporal evolution function of the vortex representation, comprising $N_{\boldsymbol{x}}(t)$, $N_{\omega}(t)$, and $N_{\delta^{2}}(t)$ as specified in Equation 4. We employ image-based supervision through the following procedure to optimize the vortex representation $V_{t}=T(t)$ at time $t$. Initially, we extract a sequence of $m+1$ consecutive frames $\left[I_{t},\ldots,I_{t+m}\right]$ from the input video. Subsequently, we derive the velocity field $\boldsymbol{u}_{t}(\boldsymbol{x})=B\left(V_{t}\right)$ using Equation 3, where $B$ represents the Biot-Savart law function incorporating $N_{K}\left(\frac{\left|\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{\prime}\right|}{\delta_{i}}\right)$ as defined in Equation 4. The pair $\left(u_{t},I_{t}\right)$ is then processed through an Eulerian grid integrator to generate the predicted frame $\tilde{I}_{t+1}$. Concurrently, $\left(\boldsymbol{u}_{t},V_{t}\right)$ undergoes integration on Lagrangian particles to yield $\tilde{V}_{t+1}$. This process iterates recursively, utilizing $\tilde{I}_{t+1}$ and $\tilde{V}_{t+1}$ as inputs to generate $\tilde{I}_{t+2}$ and $\tilde{V}_{t+2}$, and so forth. The iteration culminates in a sequence of predicted frames $\left[\tilde{I}_{t+1},\ldots,\tilde{I}_{t+m}\right]$ originating from time $t$. The optimization of $T$ and $B$ is achieved through end-to-end training by minimizing the discrepancy between the predicted sequence $\left[\tilde{I}_{t+1},\ldots,\tilde{I}_{t+m}\right]$ and the ground truth sequence $\left[I_{t+1},\ldots,I_{t+m}\right]$.
Furthermore, we employ a physical model to serve dual purposes: constraining the training process of the temporal evolution function $T$ for vortex representation and predicting the subsequent evolution of vortex flow dynamics. In particular, we establish the differential relationships governing the vortex attributes:
$\left\{\begin{array}[]{l}\frac{dN_{\boldsymbol{x}}(t)}{dt}=\boldsymbol{u}(\boldsymbol{x})\\ \frac{dN_{\omega}(t)}{dt}=\nu\nabla^{2}\omega+\nabla\times\boldsymbol{b}\\ \frac{dN_{\delta^{2}}(t)}{dt}=4\nu\end{array}\right.,$ (5)
where the temporal derivative $\frac{dN_{\boldsymbol{x}}(t)}{dt}$ is derived from the fundamental differential relationship between displacement and velocity. The expression $\frac{dN_{\omega}(t)}{dt}$ is formulated based on the vorticity transport equation, as specified in Equation 1, particularly within the framework of two-dimensional flow dynamics. Moreover, the term $\frac{dN_{\delta^{2}}(t)}{dt}$ is predicated on the decay of vortices due to viscosity, as well as the empirical processes characteristic of the Lamb-Oseen vortex [17]. Further implementation details are elaborated in Appendix A.
5 Experiments
In this study, we employ two baseline models for comparison: the PINN-based approach introduced by Raissi et al. [8], and the CNN-based methodology developed by Zhang et al. [15]. We maintain consistent training protocols to ensure a fair comparison across all models. Specifically, rather than using concentration variables as training labels in the PINN-based model, we utilize image RGB values as one of the training labels identical to our proposed model and the CNN-based approach. For the PINN-based method, we assign a weight of $10^{6}$ to the loss function component associated with image RGB values while maintaining unity weights (1.0) for all other equation-based loss function components. All images in this study maintain a consistent resolution of 512 $\times$ 512 pixels.
5.1 Single Lamb-Oseen vortex
We selected the well-established Lamb-Oseen vortex to validate our framework. The primary advantage of utilizing the Lamb-Oseen vortex lies in its analytical solution, which serves as a reliable ground truth. This vortex model characterizes the viscous decay of a vortical flow field. The two-dimensional Lamb-Oseen vortex solution is derived from the NS equations in polar coordinates $(r,\theta)$ (representing the radial and tangential directions, respectively), with velocity components $(v_{r},v_{\theta})$ expressed as:
$\left\{\begin{array}[]{l}v_{r}=0\\ v_{\theta}=\frac{\Gamma}{2\pi r}g(r,t)\\ g(r,t)=1-\mathrm{e}^{-r^{2}/4\nu t}\end{array}\right.,$ (6)
where $r$ denotes the radial distance from the vortex core and $\Gamma$ represents the circulation strength of the Lamb-Oseen vortex. As demonstrated in Equation 6, the Lamb-Oseen vortex experiences temporal decay due to viscous effects, resulting in a progressive diminution of its strength.
To validate our framework rigorously, we employed the Lamb-Oseen vortex at three distinct Reynolds numbers for training and testing datasets. The Reynolds number, a dimensionless parameter, characterizes the ratio between inertial and viscous forces within a flow. Flows with identical dimensionless parameters exhibit similar characteristics. For the Lamb-Oseen vortex, the Reynolds number (Re) is defined as:
$Re=\frac{|\Gamma|}{2\pi\nu}.$ (7)
We employ a passive scalar $s$ to visualize the vortex flows. This passive scalar $s$ satisfies the advection equation presented below:
$\frac{\partial s}{\partial t}+\boldsymbol{u}\cdot\nabla s=0.$ (8)
We establish a square computational domain with dimensions $x\in[-\pi,\pi]$ and $y\in[-\pi,\pi]$, with the Lamb-Oseen vortex core situated at the coordinate origin. The computational domain is discretized using 1024 $\times$ 1024 grid points. According to Equation 6, as $t$ approaches zero, the velocity near the vortex center becomes singular, while the vortex radius diminishes to zero. These characteristics compromise numerical accuracy and represent physical conditions rarely encountered in real-world vortex flows. Consequently, we focus our analysis on the Lamb-Oseen vortex flow at $t=1$, where the flow exhibits more physically realistic behavior. The passive scalar field is initialized as $s(x,y,t=1)=\sin(x)\sin(y)$. Throughout all simulations, we maintain a constant circulation parameter of $\Gamma=20\pi$ while systematically varying the kinematic viscosity as $\nu=1.0,0.1,0.01\text{ m}^{2}/\text{s}$ to achieve Reynolds numbers of $Re=10,100,1000$, respectively. Data is sampled at intervals of 0.01 seconds, with each sample constituting a frame in our dataset. Open boundary conditions are implemented on all boundaries of the computational domain.
Figures 3, 4, and 5 illustrate the comparative performance of the proposed method against baseline models in predicting Lamb-Oseen vortex dynamics at Reynolds numbers $Re=10$, $Re=100$, and $Re=1000$, respectively. The results demonstrate that our method achieves exceptional accuracy in hindcasting and forecasting scenarios compared to the ground truth data across all examined Reynolds numbers. Despite some discrepancies between our results and the ground truth at $Re=1000$ in the vortex core regime, we attribute these differences primarily to discretization errors in Equation 8. In contrast, while the PINN approach successfully captures most of the flow regime with reasonable fidelity, it exhibits non-physical patterns in certain regions. The CNN-based predictions, however, manifest significant numerical diffusion and rapidly deteriorate into non-physical flow characteristics as the simulation progresses.
Figure 3: Comparative evaluation of Lamb-Oseen vortex prediction performance at $Re=10$ among the proposed method and baseline models.
Figure 4: Comparative evaluation of Lamb-Oseen vortex prediction performance at $Re=100$ among the proposed method and baseline models.
Figure 5: Comparative evaluation of Lamb-Oseen vortex prediction performance at $Re=1000$ among the proposed method and baseline models.
Figure 6 presents a comprehensive quantitative evaluation of prediction accuracy for the Lamb-Oseen vortical flow field. The analysis encompasses frames 0-66, representing the observed historical data sequence, followed by frames depicting the predicted spatiotemporal evolution of flow dynamics. Six distinct performance metrics are evaluated: velocity average end-point error (AEPE), compressibility error, velocity average angular error (AAE), vorticity error, image VGG perceptual error, and image RMSE error, all assessed over time. Our proposed method demonstrates superior performance across nearly all metrics compared to the baseline approaches. The dynamics of an isolated vortex flow subjected to the combined influence of viscous effects and non-conservative body forces are examined in Appendix B.
Figure 6: Quantitative assessment of prediction accuracy for the Lamb-Oseen vortical flow field, wherein frames 0-66 comprise the observed historical data sequence, and subsequent frames represent the predicted spatiotemporal evolution of the flow dynamics.
5.2 Multiple Lamb-Oseen vortices
To evaluate our method’s capability in predicting multiple-vortex systems, we assessed its performance using a configuration comprising three distinct Lamb-Oseen vortices. Since analytical solutions for multiple Lamb-Oseen vortex systems do not exist, we employed computational fluid dynamics (CFD) methods for analysis. Specifically, we utilized the open-source CFD software OpenFOAM with the pimpleFoam solver. The computational domain consisted of a square region with dimensions $x\in[0,2\pi]$ and $y\in[0,2\pi]$, discretized using a high-resolution grid of 1024 $\times$ 1024 points. A passive scalar $s$ was initialized at $t=1$ according to $s(x,y,t=1)=\sin(2x)\sin(2y)$. The fluid viscosity was set to $\nu=0.001\text{ m}^{2}/\text{s}$. The characteristic parameters of the three Lamb-Oseen vortices are presented in Table 1.
Table 1: Parameters of Lamb-Oseen vortices in the multiple-vortex system
Index Circulation Reynolds Number Location
1 $2\pi$ 1000 $(\pi/2,\pi/2)$
2 $1.6\pi$ 800 $(3\pi/2,3\pi/2)$
3 $-1.2\pi$ 600 $(\pi,\pi)$
Figure 7 presents a comprehensive comparative analysis of multiple Lamb-Oseen vortex system predictions across the proposed methodology and established baseline models. The results indicate that the PINN-based approach accurately predicts flow characteristics in approximately half of the domain; however, a distinct region exhibits substantial inaccuracies. This discrepancy may be attributed to the increased training complexity introduced by incorporating advection equations and image RGB loss functions into the original PINN framework. The training loss convergence profile of the PINN is presented in Appendix C. The modified architecture necessitates a delicate balance between optimizing the physical equation constraints and minimizing the image RGB loss components. While the CNN-based model successfully approximates the general vortex morphology, it shows notable deficiencies in precisely resolving the vortex core regions and maintaining flow field smoothness. Although our method demonstrates similar limitations in vortex core prediction, it performs better in maintaining flow field continuity and smoothness.
Figure 7: Comprehensive comparative analysis of predictive capabilities for multiple Lamb-Oseen vortex systems among the proposed methodology and established baseline models.
Figure 8 presents a quantitative assessment of prediction accuracy for the multiple Lamb-Oseen vortical flow field. In this analysis, frames 0-66 constitute the observed historical data sequence, while subsequent frames represent the predicted spatiotemporal evolution of the flow dynamics. The results indicate that all three methods demonstrate certain advantages across various performance metrics, though none consistently outperforms the others. Therefore, our proposed method performs similarly to the baseline models in simulating multiple vortice flows. Our model’s suboptimal performance compared to the ground truth model can be primarily attributed to the absence of vortex interaction physics in our formulation.
Figure 8: Quantitative assessment of prediction accuracy for the multiple Lamb-Oseen vortical flow field, wherein frames 0-66 comprise the observed historical data sequence, and subsequent frames represent the predicted spatiotemporal evolution of the flow dynamics.
6 Conclusion
In this study, we extend the capabilities of the DVPM to non-ideal conditions, encompassing viscous flows and non-conservative body forces. Furthermore, we employ the Lamb-Oseen vortex as a more rigorous benchmark for model evaluation, as it provides an exact solution to the NS equations. However, our current framework exhibits limitations in accurately predicting multiple-vortex systems, primarily due to insufficient incorporation of the physical laws governing vortex interactions. Beyond this limitation, future work will address additional challenges, including flow transition, turbulent, and three-dimensional flow configurations.
Appendix A Implementation detail
In this section, we describe the implementation details of our proposed method.
The networks $N_{\boldsymbol{x}}$, $N_{\omega}$, and $N_{\delta^{2}}$ are implemented through a sequence of 3 residual blocks with progressively increasing widths of $[64,128,256]$. The network $N_{K}$ comprises four fully connected layers, each containing 40 neurons. Additional hyperparameters align with those established in previous research [2]. Prior to the primary training phase, $N_{\boldsymbol{x}}$ undergoes a pretraining process for 10,000 iterations with two distinct objectives: (1) ensuring that for all $t\in\left[0,t_{E}\right]$, $N_{\boldsymbol{x}}(t)=\left[\left(\boldsymbol{x}_{1}\right)_{t},\ldots,\left(\boldsymbol{x}_{n}\right)_{t}\right]$ precisely corresponds to the centers of an $n\times n$ grid, and (2) establishing that for all $t\in\left[0,t_{E}\right]$, $\frac{dN_{\boldsymbol{x}}}{dt}=0$, thereby initializing the particles in a stationary state.
Computational performance. Running on a laptop with Nvidia RTX 3090 Ti and AMD EPYC 7R12, our model takes around 0.12 s per training iteration, and around 40000 iterations to converge (for a $256\times 256$ video with 100 frames). For inference, each advance step costs around 0.01 s.
Appendix B Single Lamb-Oseen vortex with Coriolis force
[…]
Given that a Lamb-Oseen vortex under Coriolis force lacks an exact analytical solution, we employ the CFD software OpenFOAM for numerical simulation. We utilize a circular computational domain for this investigation to mitigate boundary condition effects on both internal and external flows. Figure 9 illustrates the computational domain configuration and spatial discretization methodology implemented in our CFD simulation. We partition the computational domain into five distinct blocks to enhance mesh quality. The grid point distribution for the upper-left block is explicitly indicated in the figure, while the remaining blocks maintain symmetrical grid point distributions concerning the coordinate origin. The sampling interval and boundary conditions employed in this analysis are identical to those described in Section 5.2.
Figure 9: Computational domain configuration and spatial discretizations for the CFD simulation.
Figure LABEL:lamb_oseen_vortex_Re_1000_comparison illustrates the comparative evolution of the Lamb-Oseen vortex at $t=1.2\text{s}$ ($Re=1000$) with and without the influence of the Coriolis force. The results demonstrate that when the rotational direction of the computational domain aligns with that of the Lamb-Oseen vortex (counter-clockwise rotation), the Coriolis force enhances the vortex structure. Conversely, the vortex is weakened when the rotational directions oppose each other. This observed phenomenon is consistent with the theoretical predictions derived from Equation 1.
Figure 11 presents a comparative analysis of the prediction results obtained from our proposed model and several baseline models. The results demonstrate that our model accurately predicts the single vortex flow dynamics under viscous effects and Coriolis forces. While the CNN-based model successfully captures most of the vortex flow characteristics, it exhibits notable limitations in accurately resolving the core region. In contrast, the PINN-based model yields significantly inferior results compared to the other two approaches, with substantial deviations from the expected flow patterns.
Figure 11: Rigorous comparative assessment of the predictive performance of the proposed methodology against established baseline models for Lamb-Oseen vortex under non-conservative body forces, specifically Coriolis force effects.
Appendix C Training loss curve of PINN for modeling multiple Lamb–Oseen vortices
Figure 12 illustrates the training loss curve of the PINN for modeling multiple Lamb–Oseen vortices, as described in Section 5.2. The results indicate that the loss function has reached a plateau, and further reduction appears unattainable. Despite this convergence in the training process, the results presented in Figure 7 remain unsatisfactory.
Figure 12: Training loss curve of PINN for modeling multiple Lamb–Oseen vortices.